OpenStudy (anonymous):

add simplify if possible 5/9z^3 +1/15x^2

OpenStudy (anonymous):

is the problem as followed \[\frac{5}{9z^{3}}+\frac{1}{15x^{3}}\]

jhonyy9 (jhonyy9):

(25z3 +3x2)/45

OpenStudy (anonymous):

no the last part of that is 15x2 not 3

OpenStudy (anonymous):

oops sorry, x^2

OpenStudy (anonymous):

its ok

jhonyy9 (jhonyy9):

x and z are in denominators ?

OpenStudy (anonymous):

\[\frac{25x^{2}+3z^{3}}{45z^{3}x^{2}}\]

OpenStudy (anonymous):

you first cross multiply the two fractions in order to get a common denominator then add the two fractions together and then pull out a factor of 3.

OpenStudy (anonymous):

so my answer then would be 625x+27/45

jhonyy9 (jhonyy9):

your answer is without z ???

OpenStudy (anonymous):

No, let me type it all out for you. \[\frac{5}{9z^{3}} +\frac{1}{15x^{2}}\] cross multiplying the fractions you get \[\frac{5(15x^{2})}{9z^{3}15x^{2}} +\frac{9z^{3}}{15x^{2}9z^{3}}\] which can be simplied to \[\frac{75x^{2}+9z^{3}}{135x^{2}z^{3}}\] which can then be simplifed to \[\frac{25x^{2}+3z^{3}}{45x^{2}z^{3}}\]